OpenStudy (anonymous):

Determine the first and second derivative of sin^2t-2cos^2t.

OpenStudy (anonymous):

My answer : First derivative: 2sintcost+4costsint second derivative: -2costsint+4sintcost

OpenStudy (anonymous):

But it's wrong.

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (loser66):

do you know why it's wrong? It's not wrong, just not the simplest form

OpenStudy (loser66):

add them together, you have 6sintcost, right? for the first derivative

OpenStudy (loser66):

and it is = 3(2 sint cos t) right?

OpenStudy (loser66):

= 3 sin(2t) . that is the simplest form

OpenStudy (loser66):

then, take the second derivative

OpenStudy (anonymous):

Second derivative is incorrect. Use the product rule. Alternatively, you may wish to use identities for the given function to condense the work: \[\sin^2t-2\cos^2t=\frac{1-\cos2t}{2}-2\left(\frac{1+\cos2t}{2}\right)=-\frac{1}{2}-\frac{3}{2}\cos2t\] Then the first derivative is \[3\sin2t=3(2\sin t\cos t)=6\sin t\cos t\] and the second derivative is \[6\cos2t=6\cos^2t-6\sin^2t=\text{whatever else this may be equivalent to}\]